OpenStudy (anonymous):

find the area of the surface generated by revolving the curve: x=t, y=2t, 0

OpenStudy (anonymous):

So making a cone with its tip at z,y,z = 0 and open end in the positive x direction. Should be \[2\pi * magnitude\] where magnitude is the length of the line in the x-y plane. It is a straight line from 0,0 to 3,6, magnitude is \[\sqrt{3^2+6^2} = \sqrt{45}\] so the area around the cone is \[2*\pi*\sqrt{45}\]